40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2080     Entry time: Thu Feb 8 16:48:07 2018
Author: awade, Craig 
Type: DailyProgress 
Category: NoiseBudget 
Subject: BN spectrum Feb 08, 2018 

Checking PLL OLG

Measured the OLG using out1/out2 method.  Did the G/(1+G) invertion to OLG directly on the SR785. Plotted in attachement 1.

Looks like a UGF of about 16.5 kHz

BN spectrum

Current beat note is at 8.98 MHz.  North cavity heater is set to 0.77422 W and will probably need to be raised a little to get us a little higher in BN frequency.

PLL settings are:
Marconi FM Devn: 10 kHz
SR560 Gain: 20
Beat note strength: -1.17 dBm

Calibrated BN spectrum is attached below, attachment 2

 

 

 

Attachment 1: 20180208_172015_PLLOLTF_08-02-2018_171734.pdf  97 kB  | Hide | Hide all
20180208_172015_PLLOLTF_08-02-2018_171734.pdf
Attachment 2: 20180208_175847_CalibratedSpectrum.pdf  145 kB  Uploaded Thu Feb 8 18:19:15 2018  | Hide | Hide all
20180208_175847_CalibratedSpectrum.pdf
Attachment 3: 20180208_dailyBNSpectrum.tar.gz  95 kB  Uploaded Thu Feb 8 19:33:48 2018
ELOG V3.1.3-